# A Calculus of Inheritance

### 10 Mar 2026 v6

@@ 7.3. Configuration Languages and Module Systems
CUE (van Lohuizen, 2019) unifies types and values in a single lattice where unification (&) is commutative, associative, and idempotent. CUE’s design was motivated by Google’s Borg Configuration Language (GCL), which used graph unification. CUE’s lattice includes scalar types with a conflict semantics where unifying incompatible scalars yields $\bot$; whether scalars are necessary for a configuration language, or whether tree structure alone suffices, is an interesting design question. Jsonnet (Cunningham, 2014) provides object inheritance via the + operator with mixin semantics: composition is not commutative since the right-hand side wins on scalar conflicts, and deep merging requires explicit opt-in via the +: syntax) on a per-field basis, so deep merge is available but not the default. Dhall (Gonzalez, 2017) takes a functional approach: it is a typed $\lambda$-calculus with records, where composition is a record merge operator, not an inheritance mechanism.
@@
The NixOS module system (Dolstra et al., 2010; NixOS Contributors, 2024) composes modules by recursively merging nested attribute sets, not by linearization or method-level override, which is the same deep-merge semantics that distinguishes tree-level inheritance from method-level composition in prior mixin and trait calculi. When two modules define the same nested path, their subtrees are merged rather than overridden. The NixOS module system incorporates deep merge, recursive self-reference, deferred evaluation of modules, and a rich type-checking layer on top. The same mechanism powers NixOS system configuration, Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), flake-parts (Hensing, 2022), and dozens of other ecosystems, collectively managing configurations of arbitrary complexity. The one aspect that appears not to have been fully addressed is the semantics of self-reference under multiple inheritance. When two independent modules that inherit from the same deferred module are composed, each import introduces its own declaration of the shared options; the module system’s mergeOptionDecls rejects duplicate declarations with a static error (Appendix D). The module system cannot express that both routes contribute equally to the same option, that is, the multi-path situation that recursive merging naturally gives rise to, which motivated the observational semantics of Section 3.
@@ 7.4. Our Companion Work
Union file systems (UnionFS, OverlayFS, AUFS) layer multiple directory hierarchies to present a unified view. Their semantics resembles inheritance-calculus in that later layers override earlier layers and files from all layers remain accessible, but without a formal foundation these systems resort to ad-hoc choices: asymmetric inheritance, noncommutative conflict resolution for scalar files, and no late-binding or dynamic dispatch across layers. We implemented inheritance-calculus directly as a union file system (Yang, 2025), obtaining late-binding semantics and dynamic dispatch for file lookups with no additional mechanism beyond the three primitives. In package managers, build systems, and OS distributions, external toolchains transform configuration into file trees; here, the file system serves as configuration to itself, and the transformation is inheritance: a file-system-as-compiler in which the directory tree is both the source program and the compilation target.
@@ 8.1. Type System
MIXINv2’s existing compile-time checks verify that every reference path resolves to a property that exists in the inherited result. Formalizing the soundness of these checks, that is, proving that well-typed programs do not produce dangling references at runtime, is one direction of future work. Such a formalization would relate to the DOT calculus (Amin et al., 2016) in a manner analogous to how inheritance-calculus relates to the $\lambda$-calculus: a Typed inheritance-calculus can be viewed as DOT without $\lambda$, retaining path-dependent types while replacing functions with mixin inheritance. A formal type system would be amenable to mechanization in proof assistants (Coq, Agda) for verifying type safety, progress, and preservation properties.
@@ 8.2. Standard Library
Inheritance-calculus is open by default: inheritance can freely merge any two mixins, and a value may simultaneously inhabit multiple data constructors (e.g., $\{\mathrm{Zero},\;\mathrm{Odd},\;\mathrm{half}\mapsto\mathrm{Zero}\}$). This is the natural trie semantics of the calculus, and the basis for solving the expression problem (Section 5). However, many operations (e.g., equality testing) assume that values are linear, inhabiting exactly one data constructor. The observer pattern (Section 5) does not enforce this: applied to a multi-data-constructor value, all callbacks fire and their results are inherited.
@@ 9. Conclusion
Immunity to nonextensibility guided the design of inheritance-calculus. The emergent phenomena exhibit an expressive asymmetry between inheritance-calculus and the $\lambda$-calculus. Functions are syntactic sugar for inheritance.
@@ Appendix A Well-Definedness of the Semantic Functions
The six semantic equations (1)–(6) are first-order: their inputs and outputs are paths and sets of paths, with no function spaces, closures, or higher-order constructs. This section establishes well-definedness by showing that the six equations define a monotone operator whose least fixed point is computed soundly by tabled evaluation (Tamaki and Sato, 1986; Chen and Warren, 1996).
@@
    Monotone accumulation. The tabling accumulator only grows between iterations (via set union); no fact is ever removed.
@@
    Correct fixpoint detection. If no sub-query was re-entered during evaluation, all sub-queries were fully computed with no approximation, and the result is exact. If re-entry occurred but the accumulator is stable (no new facts were added), the accumulator equals $\mathrm{lfp}(T_{P})$ restricted to the reachable queries, because it is a pre-fixed point reached by ascending from $\bot$.
@@
For (3), when no reentry occurs, the evaluation is ordinary memoised recursion on a finite acyclic reachable subgraph, so every sub-query is computed exactly and no approximation is involved. When reentry occurs, let $R\subseteq\mathcal{Q}$ be the set of queries reachable from the initial query, and let $T_{P}^{R}$ denote the restriction of $T_{P}$ to maps on $R$. If the accumulator is stable, i.e. $A=A\cup T_{P}^{R}(A)$, then $T_{P}^{R}(A)\subseteq A$, so $A$ is a pre-fixed point of $T_{P}^{R}$.
@@
Therefore $A=\mathrm{lfp}(T_{P}^{R})$, i.e. the stable accumulator equals the exact denotation on the reachable queries.
@@
  • Infinite reachable subgraph. Evaluation may diverge; this is inherent to Turing-completeness and not a defect of the evaluation strategy.
@@ Appendix B Böhm Tree Correspondence: Proofs
This appendix contains the full proofs of Adequacy (Theorem 4.2) and the first-order semantics theorem (Theorem 4.3) stated in Section 4.1.
@@ B.1. Böhm Trees
We briefly recall the definition of Böhm trees (Barendregt, 1984). A head reduction $M\to_{h}M^{\prime}$ contracts the outermost $\beta$-redex only: if $M$ has the form $(\lambda x.\,e)\;v\;M_{1}\cdots M_{k}$, then $M\to_{h}e[v/x]\;M_{1}\cdots M_{k}$. A term $M$ is in head normal form (HNF) if it has no head redex, i.e., $M=\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}$ where $y$ is a variable.
@@
Böhm tree equivalence $\mathrm{BT}(M)=\mathrm{BT}(N)$ is the standard observational equivalence of the lazy $\lambda$-calculus (Abramsky and Ong, 1993).
@@ B.2. Path Encoding
  • An abstraction $\lambda x.\,M$ translates to a record with own properties $\{\mathrm{argument},\mathrm{result}\}$ (the abstraction shape).
@@
Each Böhm tree position is a sequence of navigation steps from the root. In the Böhm tree, the possible steps at a node $\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}$ are: entering the body (peeling off one abstraction layer) and entering the $i$-th argument $M_{i}$ of the head variable. In the mixin tree, these correspond to following the $\mathrm{result}$ label (for entering the body) and the $\mathrm{argument}$ label after $\mathtt{this}$ resolution (for entering an argument position).
Rather than comparing absolute paths across different mixin trees (which would be sensitive to internal structure), we observe only the convergence behavior accessible via the $\mathrm{result}$ projection. Recall the definition of inheritance-convergence (Definition 4.1). Each $\mathrm{result}$-following step corresponds to one head reduction step in the $\lambda$-calculus: a tail call’s $\mathrm{result}$ projects through the encapsulated application, and a $\mathbf{let}$-binding’s $\mathrm{result}$ enters the continuation. The abstraction shape at the end signals that a weak head normal form has been reached.
@@ B.3. Single-Path Lemma
The $\mathtt{this}$ function (equation 6) is designed for the general case of multi-path mixin inheritance. For translated $\lambda$-terms, we show it degenerates to single-path resolution.
Lemma B.1 (Single Path).
@@ B.4. Substitution Lemma
Let $M$ be an ANF term in which $x$ may occur free, and let $V$ be a closed value. Define the inherited tree $C=\{\mathcal{T}(\lambda x.\,M),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$. Then for every path $p$ under $C\mathbin{\triangleright}\mathrm{result}$ and every label $\ell$:
@@
The substitution distributes: $M[V/x]=\mathbf{let}\;z=V^{\prime}[V/x]\;\mathbf{in}\;M^{\prime}[V/x]$. By induction on $V^{\prime}$ and $M^{\prime}$, the properties under both subtrees ($z$ and $\mathrm{result}$) coincide. In particular, occurrences of $z$ in $M^{\prime}$ are translated as the lexical reference $z.\mathrm{result}$ on both sides, so the wrapper introduced by the value $\mathbf{let}$-binding is preserved by the substitution.
@@ B.5. Convergence Preservation
The convergence criterion (Definition 4.1) follows the $\mathrm{result}$ chain from the root. Each step in this chain corresponds to one head reduction step in the $\lambda$-calculus. We make this precise.
@@
Let $(\lambda x.\,M)\;V$ be a tail call (a $\beta$-redex). Then for every path $p$ and label $\ell$:
@@
Extension to $\mathrm{lfp}(T_{P})$ semantics
The proof above works directly with the six equations (1)–(6), not with any particular evaluation strategy. We distinguish two consequences of switching from single-pass evaluation to multi-pass tabled evaluation (computing $\mathrm{lfp}(T_{P})$ via iterated application of $T_{P}$; see Appendix A).
Head-reduction preservation is maintained.:
The Result-Step Lemma (Lemma B.3) establishes a structural isomorphism between the equation systems of $\mathcal{T}(M)$ and $\mathcal{T}(N)$, shifted by one $\mathrm{result}$ step. This isomorphism preserves monotonicity, so $\mathrm{lfp}(T_{P})$ yields the same answer on both sides (modulo the depth shift): if $\mathcal{T}(M)$ converges at depth $n$ under $\mathrm{lfp}$, then $\mathcal{T}(N)$ converges at depth n1n-1, and vice versa. For $\lambda$-calculus–translated programs, the Single-Path Lemma (Lemma B.1) ensures that no cycles arise in the dependency graph (each scope level has exactly one inheritance source), so single-pass and multi-pass evaluation produce identical results in this fragment.
Agreement with the $\lambda$-calculus is not maintained in general.:
For programs beyond the $\lambda$-calculus fragment that introduce cyclic dependencies, $\mathrm{lfp}(T_{P})$ may converge where the $\lambda$-calculus diverges. This means $\mathrm{lfp}(T_{P})$ evaluation is strictly more defined than $\beta$-reduction on the translated image: $M{\Downarrow}_{\lambda}$ implies $\mathcal{T}(M){\Downarrow}_{\mathrm{lfp}}$, but not conversely.
As a minimal illustration, the $\lambda$-term $\mathbf{let}\;x=x\;\mathbf{in}\;x$ diverges under $\beta$-reduction (unfolding $x$ produces $x$ again, ad infinitum). Translation $\mathcal{T}$ gives
$\{x\mapsto\{\mathrm{result}\mapsto x.\mathrm{result}\},\;\mathrm{result}\mapsto x.\mathrm{result}\}$
where $x.\mathrm{result}$ inherits from itself, creating a cyclic dependency. Under single-pass evaluation ($\mathit{max\_fixpoint\_iterations}=0$), the cycle raises $\bot$, matching the $\lambda$-calculus divergence. Under $\mathrm{lfp}(T_{P})$, fixpoint iteration detects the cycle and converges: $\mathrm{properties}(\mathrm{root})=\{x,\mathrm{result}\}$ and $\mathrm{properties}(\mathrm{root}\triangleright\mathrm{result})=\varnothing$, reflecting that the value is undefined (no abstraction shape). Appendix F gives a more motivated example (Datalog transitive closure on cyclic graphs).

### 16 Mar 2026 v7

@@ 7.3. Configuration Languages and Module Systems
CUE (van Lohuizen, 2019) unifies types and values in a single lattice where unification, denoted by &, is commutative, associative, and idempotent. CUE’s design was motivated by Google’s Borg Configuration Language (GCL), which used graph unification. CUE’s lattice includes scalar types with a conflict semantics where unifying incompatible scalars yields $\bot$; whether scalars are necessary for a configuration language, or whether tree structure alone suffices, is an interesting design question. Jsonnet (Cunningham, 2014) provides object inheritance via the + operator with mixin semantics: composition is not commutative since the right-hand side wins on scalar conflicts, and deep merging requires explicit opt-in via the +: syntax on a per-field basis, so deep merge is available but not the default. Dhall (Gonzalez, 2017) takes a functional approach: it is a typed $\lambda$-calculus with records, where composition is a record merge operator, not an inheritance mechanism.
@@
The NixOS module system (Dolstra et al., 2010; NixOS Contributors, ) composes modules by recursively merging nested attribute sets, not by linearization or method-level override, which is the same deep-merge semantics that distinguishes tree-level inheritance from method-level composition in prior mixin and trait calculi. When two modules define the same nested path, their subtrees are merged rather than overridden. The NixOS module system incorporates deep merge, recursive $\mathtt{this}$, deferred evaluation of modules, and a rich type-checking layer on top. The same mechanism powers NixOS system configuration, Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), flake-parts (Hensing, 2022), and dozens of other ecosystems, collectively managing configurations of arbitrary complexity. The one aspect that appears not to have been fully addressed is the semantics of $\mathtt{this}$ under multiple inheritance. When two independent modules that inherit from the same deferred module are composed, each import introduces its own declaration of the shared options; the module system’s mergeOptionDecls rejects duplicate declarations with a static error (Appendix D). The module system cannot express that both routes contribute equally to the same option, that is, the multi-target situation that recursive merging naturally gives rise to, which motivated the observational semantics of Section 3.
@@ 7.4. Our Companion Work
Union file systems, such as UnionFS, OverlayFS, and AUFS, layer multiple directory hierarchies to present a unified view. Their semantics resembles inheritance-calculus in that later layers override earlier layers and files from all layers remain accessible, but their designs differ in specific ways: inheritance is asymmetric, conflict resolution for scalar files is noncommutative, and there is no late-binding or dynamic dispatch across layers. We implemented inheritance-calculus directly as a union file system (Yang, 2025), obtaining late-binding semantics and dynamic dispatch for file lookups with no additional mechanism beyond the three primitives. In package managers, build systems, and OS distributions, external toolchains transform configuration into file trees; here, the file system serves as configuration to itself, and the transformation is inheritance: a file-system-as-compiler in which the directory tree is both the source program and the compilation target.
@@ 8.1. Type System
MIXINv2’s existing compile-time checks verify that every reference path resolves to a property that exists in the inherited result. Formalizing the soundness of these checks, that is, proving that well-typed programs do not produce dangling references at runtime, is one direction of future work. Such a formalization would relate to the DOT calculus (Amin et al., 2016) in a manner analogous to how inheritance-calculus relates to the $\lambda$-calculus: a Typed inheritance-calculus can be viewed as DOT without $\lambda$, retaining path-dependent types while replacing functions with mixin inheritance. A formal type system would be amenable to mechanization in proof assistants, such as Coq and Agda, for verifying type safety, progress, and preservation properties.
@@ 8.2. Standard Library
Inheritance-calculus is open by default: inheritance can freely merge any two mixins, and a value may simultaneously inhabit multiple data constructors (e.g., $\{\mathrm{Zero},\;\mathrm{Odd},\;\mathrm{half}\mapsto\mathrm{Zero}\}$). This is the natural trie semantics of the calculus, and the basis for solving the Expression Problem (Section 5). However, many operations, such as equality testing, assume that values are linear, inhabiting exactly one data constructor. The observer pattern (Section 5) does not enforce this: applied to a multi-data-constructor value, all callbacks fire and their results are inherited.
@@ 9. Conclusion
Immunity to nonextensibility guided the design of inheritance-calculus. Section 6.2 eliminated every alternative composition mechanism and reference resolution, leaving deep merge with multi-target late-binding qualified $\mathtt{this}$ as the unique survivor. The mixin served as the single abstraction unifying module, class, object, method, field, and local binding; deep merge of mixin trees made inheritance commutative, idempotent, and associative, so the linearization problem of multiple inheritance did not arise. The six first-order equations of Section 3, denotational and computable by tabling (Tamaki and Sato, 1986), suffice to define the semantics; the observer drives computation by querying paths in a lazily constructed tree. These equations are a minimal departure from standard object-oriented semantics: the only change is the $\mathtt{overrides}$ equation (3), which replaces method-level shadowing with deep merge of subtrees. Object-oriented programming is traditionally classified as imperative; removing mutable state and function bodies yields pure object-oriented programming,313131“Pure” carries a double meaning: not impure, since values are immutable and composition has no side effects; and not mixed, since composition relies entirely on inheritance, without functional abstractions such as $\lambda$ or $\beta$-reduction. a minimal computational model for declarative programming that fills the empty cell of Table 1.
Since $\mathtt{overrides}$ merges subtrees rather than method bodies, functions are no longer primitive. The $\lambda$-calculus macro-expressed into inheritance-calculus in six translation rules in Section 4. The Böhm tree correspondence of Section 4.1 extended this first-order semantics to the $\lambda$-calculus: the prior higher-order semantic frameworks proved faithful to the first-iteration approximation of a sublanguage of inheritance-calculus’s first-order equations; Appendix E reformulated this semantics directly on ANF. A convergence hierarchy among the eager, lazy, and fixpoint $\lambda$-calculi sharing the same $\lambda$-syntax followed as a corollary, detailed in Remark B.5.
Inheritance-calculus further gains three phenomena that the $\lambda$-calculus cannot express; indeed, inheritance-calculus is strictly more expressive than the lazy $\lambda$-calculus in Felleisen’s sense (Felleisen, 1991) (Section 4.2). Section 5 established immunity to nonextensibility in the sense of the Expression Problem (Wadler, 1998). The case study of Section 5 showed that ordinary arithmetic yields the relational semantics of logic programming: cyclic inheritance hierarchies converge by the same ascending chain condition that governs recursive Datalog, and Appendix F gave a systematic encoding. Section Function color blindness: exhibited function color blindness (Nystrom, 2015).
Inheritance is not a mechanism for organizing functions. Functions are syntactic sugar for inheritance.
@@ Appendix A Well-Definedness of the Semantic Functions
The six semantic equations (1)–(6) are first-order: their inputs and outputs are paths and sets of paths, with no function spaces, closures, or higher-order constructs. This section establishes well-definedness by showing that the six equations define a monotone operator whose least fixed point is computed soundly by tabled evaluation (Tamaki and Sato, 1986).
@@
    Monotone accumulation. The tabling accumulator only grows between iterations via set union; no fact is ever removed.
@@
    Correct fixpoint detection. If no sub-query was re-entered during evaluation, all sub-queries were fully computed with no approximation, and the result is exact. If re-entry occurred but the accumulator is stable, meaning no new facts were added, the accumulator equals $\mathrm{lfp}(T_{P})$ restricted to the reachable queries, because it is a pre-fixed point reached by ascending from $\bot$.
@@
For (3), when no reentry occurs, the evaluation is ordinary memoised recursion on a finite acyclic reachable subgraph, so every sub-query is computed exactly and no approximation is involved. When reentry occurs, let $R\subseteq\mathcal{Q}$ be the set of queries reachable from the initial query, and let $T_{P}^{R}$ denote the restriction of $T_{P}$ to maps on $R$. If the accumulator is stable, that is, $A=A\cup T_{P}^{R}(A)$, then $T_{P}^{R}(A)\subseteq A$, so $A$ is a pre-fixed point of $T_{P}^{R}$.
@@
Therefore $A=\mathrm{lfp}(T_{P}^{R})$, that is, the stable accumulator equals the exact denotation on the reachable queries.
@@
  • Infinite reachable subgraph. Evaluation may diverge; this is inherent to Turing completeness and not a defect of the evaluation strategy.
@@ Appendix B Böhm Tree Correspondence: Proofs
This appendix proves two results about the sublanguage of inheritance-calculus corresponding to the $\lambda$-calculus (Section 4.1). First, the first-iteration approximation $T_{P}\!\uparrow\!1$ is fully abstract with respect to Böhm tree equivalence (Theorems B.7 and B.9), establishing that the first iteration coincides with the lazy $\lambda$-calculus. Second, the full $\mathrm{lfp}(T_{P})$ semantics, the fixpoint $\lambda$-calculus, is strictly more defined: $\lambda$-terms that diverge under $\beta$-reduction may converge (Remark B.5), establishing a convergence hierarchy among the eager, lazy, and fixpoint $\lambda$-calculi.
@@ B.1. Böhm Trees
We briefly recall the definition of Böhm trees (Barendregt, 1984). A head reduction $M\to_{h}M^{\prime}$ contracts the outermost $\beta$-redex only: if $M$ has the form $(\lambda x.\,e)\;v\;M_{1}\cdots M_{k}$, then $M\to_{h}e[v/x]\;M_{1}\cdots M_{k}$. A term $M$ is in head normal form (HNF) if it has no head redex, that is, $M=\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}$ where $y$ is a variable.
@@
Böhm tree equivalence $\mathrm{BT}(M)=\mathrm{BT}(N)$ is the standard observational equivalence of the lazy $\lambda$-calculus (Abramsky and Ong, 1993; Abramsky and McCusker, 1995).
@@ B.2. Path Encoding
  • An abstraction $\lambda x.\,M$ translates to a record with own properties $\{\mathrm{argument},\mathrm{result}\}$, which is referred to as the abstraction shape.
@@
Each Böhm tree position is a sequence of navigation steps from the root. In the Böhm tree, the possible steps at a node $\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}$ are: entering the body by peeling off one abstraction layer, and entering the $i$-th argument $M_{i}$ of the head variable. In the mixin tree, these correspond to following the $\mathrm{result}$ label (for entering the body) and the $\mathrm{argument}$ label after $\mathtt{this}$ resolution (for entering an argument position).
Rather than comparing absolute paths across different mixin trees, which would be sensitive to internal structure; hence, we observe only the convergence behavior accessible via the $\mathrm{result}$ projection. Recall the definition of inheritance-convergence (Definition B.6). Each $\mathrm{result}$-following step corresponds to one head reduction step in the $\lambda$-calculus: a tail call’s $\mathrm{result}$ projects through the encapsulated application, and a $\mathbf{let}$-binding’s $\mathrm{result}$ enters the continuation. The abstraction shape at the end signals that a head normal form has been reached.
@@ B.3. Single-Path Lemma
The $\mathtt{this}$ function (equation 6) is designed for the general case of multi-target $\mathtt{this}$ resolution. For translated $\lambda$-terms, we show it degenerates to single-target resolution.
Lemma B.1 (Single-Path).
@@ B.4. Substitution Lemma
Let $M$ be an ANF term in which $x$ may occur free, and let $V$ be a closed value. [^32] The proof uses the Single-Path Lemma (Lemma B.1), which requires the inherited tree to occur within a closed term. This holds whenever $\lambda x.\,M$ applied to $V$ is a subterm of a closed ANF program. Define the inherited tree $C=\{\mathcal{T}(\lambda x.\,M),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$. Then for every path $p$ under $C\mathbin{\triangleright}\mathrm{result}$ and every label $\ell$:
@@
The substitution distributes: $M[V/x]=\mathbf{let}\;z=V^{\prime}[V/x]\;\mathbf{in}\;M^{\prime}[V/x]$. By induction on $V^{\prime}$ and $M^{\prime}$, the properties under both subtrees, specifically $z$ and $\mathrm{result}$, coincide. In particular, occurrences of $z$ in $M^{\prime}$ are translated as the lexical reference $z.\mathrm{result}$ on both sides, so the wrapper introduced by the value $\mathbf{let}$-binding is preserved by the substitution.
@@ B.5. Convergence Preservation
The convergence criterion (Definition B.6) follows the $\mathrm{result}$ chain from the root. Each step in this chain corresponds to one head reduction step in the $\lambda$-calculus. We make this precise.
@@
Let $(\lambda x.\,M)\;V$ be a tail call, which is a $\beta$-redex. Then for every path $p$ and label $\ell$:
@@
Remark B.5 (Convergence hierarchy).
Section 4.1 showed that the Böhm tree correspondence holds at the first iteration $T_{P}\!\uparrow\!1$ and that $\mathrm{lfp}(T_{P})$ is strictly more defined. Here we make the connection precise. Tabled evaluation (Tamaki and Sato, 1986) computes $\mathrm{lfp}(T_{P})$ by iterating: $T_{P}\!\uparrow\!0=\bot$, $T_{P}\!\uparrow\!1=T_{P}(\bot)$, $T_{P}\!\uparrow\!2=T_{P}(T_{P}(\bot))$, … until stabilisation (Theorem A.6).
$T_{P}\!\uparrow\!1$ corresponds to $\beta$-reduction.:
$T_{P}\!\uparrow\!1$ evaluates each equation once, with re-entrant calls returning $\bot$ (the value from $T_{P}\!\uparrow\!0$). The Result-Step Lemma (Lemma B.3) establishes a structural isomorphism between the equation systems of $\mathcal{T}(M)$ and $\mathcal{T}(N)$, shifted by one $\mathrm{result}$ step. This isomorphism is purely structural and holds for any number of iterations. Theorems B.7 and B.9 are proved at this level.
$\mathrm{lfp}(T_{P})$ defines the fixpoint $\lambda$-calculus.:
We call the sublanguage with $\mathrm{lfp}(T_{P})$ semantics the fixpoint $\lambda$-calculus; it shares the same syntax as the lazy $\lambda$-calculus but has strictly greater computational power: $M{\Downarrow}_{\lambda}$ implies $\mathcal{T}(M){\Downarrow}_{\mathrm{lfp}}$, but not conversely.